UNIVERSITY of CALIFORNIA

SANTA CRUZ

THE ELEMENTS OF THESES AS A DEMONSTRATION OF

UCSCTHESIS.CLS FOR SECOND EDITION L^{A}TEX

A thesis submitted in partial satisfaction of the requirements for the degree of BACHELOR OF SCIENCE

in PHYSICS by Perry H. Disdainful 10 June 2008

The thesis of Perry H. Disdainful is approved by:

Professor Absent Minded Theses Coordinator

Professor Ivory Insular Advisor

Professor David P. Belanger Chair, Department of Physics

Copyright \begin{matrix}c \\ ◯ \end{matrix} by Perry H. Disdainful

Abstract

The Elements of Theses as a Demonstration of ucscthesis.cls for Second Edition L^{A}TEX

by Perry H. Disdainful

Theses have elements. Isn’t that nice? All work and no play makes Johnny a dull boy. All work and no play makes Johnny a dull boy. All work and no play makes Johnny a dull boy. All work and no play makes Johnny a dull boy. All work and no play makes Johnny a dull boy. All work and no play makes Johnny a dull boy. All work and no play makes Johnny a dull boy. All work and no play makes Johnny a dull boy. All work and no play makes Johnny a dull boy. All work and no play makes Johnny a dull boy. All work and no play makes Johnny a dull boy. All work and no play makes Johnny a dull boy. All work and no play makes Johnny a dull boy. All work and no play makes Johnny a dull boy.

Contents

List of Figures v List of Tables vi Dedication vii Acknowledgements viii 1 Introduction 1

1.1 This demonstrates a section . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1.1 Subsection 1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2 Previous Work 4 3 Conclusion 6 A Some Ancillary Stuff 7 Bibliography 8

List of Figures

1.1 Transverse Scans at difference Temperatures at H = 11 T . . . . . . . . . . . . . . . 2 2.1 A first figure. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.2 A second figure. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

List of Tables

1.1 A normalsize table. There has been a complaint that table captions are not single-

spaced. This is odd because the code indicates that they should be. . . . . . . . . . 3 1.2 A small table. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

To myself, Perry H. Disdainful,

the only person worthy of my company.

Acknowledgements

I want to “thank” my committee, without whose ridiculous demands, I would have graduated so, so, very much faster.

1 Introduction

Every thesis should have an introduction. You might not realize it, but the introduction

should introduce the concepts, backgrouand, and goals of the thesis. Every thesis should have an introduction. You might not realize Every thesis should have an introduction. You might not realize Every thesis should have an introduction. You might not realize it, but the introduction should introduce the concepts, backgrouand, and goals of the thesis. it, but the introduction should introduce the concepts, backgrouand, and goals of the thesis. it, but the introduction should introduce the concepts, backgrouand, and goals of the thesis.

1.1 This demonstrates a section

Here’s some text for the section. Here’s some text for the section. Here’s some text for the

section. Here’s some text for the section. Here’s some text for the section. Here’s some text for the section.

Here’s some text for the section. Here’s some text for the section. Here’s some Fig. 1.1

text for the section. Here’s some text for the section. Here’s some text for the section. Here’s some text for the section. Here’s some text for the section. Here’s some text for the section. Here’s some text for the section. Here’s some text for the section. Here’s some text for the section. Here’s some text ν^{2β} for the section. Here’s some text for the section. Here’s some text for the section.

Figure 1.1: Representative transverse scans for the different temperatures below T_{c}(H) = 63.7 K taken with H = 11 T after cooling in H = 0. Each scan is displaced vertically by 0.1 units from the scan below it for clarity. The solid curves are results of least-squares fits to a Gaussian line shapes with a half-width-at-half-maximum equal to 2.1 × 10^{−4} r.l.u.

1.1.1 Subsection 1

Here’s a citation of one of the references [1]. Here’s some text to fill up the sub-section.

Here’s some text to fill up the sub-section. Here’s some text to fill up the sub-section. We can do equations like this first famous attempt at special relativity E = ma^{2}. A second attempt was E = mb^{2}. Finally, the correct answer was found to be

E = mc^{2}  (1.1)

However, a pathetic attempt was made to push for

E = mc_{2}  (1.2)

but that was quickly rejected with much ridicule. Here’s some text to fill up the sub-section. Here’s some text to fill up the sub-section. Here’s some text to fill up the sub-section. Here’s some text to

Title  Author
War And Peace  Leo Tolstoy
The Great Gatsby F. Scott Fitzgerald

Table 1.1: A normalsize table. There has been a complaint that table captions are not single-spaced. This is odd because the code indicates that they should be.

Table 1.2: A small table.

Title  Author
War And Peace  Leo Tolstoy
The Great Gatsby F. Scott Fitzgerald

fill up the sub-section. Here’s some text to fill up the sub-section. Here’s some text to fill up the sub-section. Here’s some text to fill up the sub-section. Here’s some text to fill up the sub-section. Here’s some text to fill up the sub-section.

Here’s some text to fill up the sub-section. Here’s some text to fill up the sub-section.

Here’s some text to fill up the sub-section. Here’s some text to fill up the sub-section. Here’s some text to fill up the sub-section. Here’s some text to fill up the sub-section. Here’s some text to fill up the sub-section.

2 Previous Work

Some other research was once performed. Some other research was once performed. Some

other research was once performed. Some other research was once performed. Some other research was once performed. Some other research was once performed. Some other research was once performed. Some other research was once performed. Some other research was once performed. Some other research was once performed. Some other research was once performed.

Some other research was once performed. Some other research was once performed. Some

other research was once performed. Some other research was once performed. Some other research was once performed. Some other research was once performed. Some other research was once performed. Some other research was once performed.

Some other research was once performed. Some other research was once performed. Some

other research was once performed. Some other research was once performed. Some other research was once performed. Some other research was once performed. Some other research was once performed. Some other research was once performed.

Figure 2.1: A first figure.

5

Figure 2.2: A second figure.

3 Conclusion

This is where a conclusion would go. This is where a conclusion would go. This is where

a conclusion would go. This is where a conclusion would go. This is where a conclusion would go. This is where a conclusion would go. This is where a conclusion would go. This is where a conclusion would go. This is where a conclusion would go. This is where a conclusion would go. This is where a conclusion would go.

This is where a conclusion would go. This is where a conclusion would go. This is where

a conclusion would go. This is where a conclusion would go. This is where a conclusion would go. This is where a conclusion would go. This is where a conclusion would go. This is where a conclusion would go.

This is where a conclusion would go. This is where a conclusion would go. This is where

a conclusion would go. This is where a conclusion would go. This is where a conclusion would go. This is where a conclusion would go. This is where a conclusion would go. This is where a conclusion would go.

Appendix A

Some Ancillary Stuff

Ancillary material should be put in appendices. The guidelines are not clear whether

bibliography comes before or after the appendices, but they suggest appendices come first. Ancillary material should be put in appendices. The guidelines are not clear whether bibliography comes before or after the appendices, but they suggest appendices come first. Ancillary material should be put in appendices. The guidelines are not clear whether bibliography comes before or after the appendices, but they suggest appendices come first. Ancillary material should be put in appendices. The guidelines are not clear whether bibliography comes before or after the appendices, but they suggest appendices come first.

Bibliography

[1] E. T. Seppälä, A. M. Pulkkinen, and M. J. Alava, Phys. Rev. B 66, 144403 (2002).